# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@ 3.2. Proof of Theorem A (ii)
We prove the statement by contradiction. Assume that there exists $x_{0}\in X$ such that $f(x_{0})\neq g(x_{0})$. Since $Y$ is Hausdorff, there exist open neighborhoods $U$ and $V$ of $f(x_{0})$ and $g(x_{0})$, respectively, with $U\cap V=\varnothing$. Observe now that $W=f^{-1}(U)\cap g^{-1}(V)\subset X$ is a non-empty open subset. Since $D$ is dense in $X$, there exists $y\in D\cap W$ and, by hypothesis, $f(y)=g(y)$. Since $y\in W$, $f(y)\in U$ and $g(y)\in V$. This implies that $V\cap U\neq\varnothing$, which is a contradiction. ∎
Proof of Theorem A (ii).
Recall that $M_{\mathrm{prin}}\subset M$, the set of principal leaves, is an open, connected, and dense subset of $M$ (see Section 2.1). The quotient $M_{\mathrm{prin}}/\mathcal{F}$ is a Riemannian manifold. Hence, by the Myers–Steenrod theorem, $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is a Lie group. Moreover, $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is closed (for example, by the proof of Lemma 3.4). In addition, the projection map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is proper. These observations, along with Proposition 3.1, imply that $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ is a closed subgroup of $\operatorname{Isom}(M_{\mathrm{prin}})$. Now, since $M_{\mathrm{prin}}$ is a smooth submanifold of $M$, $\operatorname{Isom}(M_{\mathrm{prin}})$ is a Lie group. Therefore, $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ is a Lie group.
Denote the identity map on $M$ by $\mathrm{Id}_{M}$ and its restriction to $M_{\mathrm{prin}}$ by $\mathrm{Id}_{M_{\mathrm{prin}}}$. Consider the homomorphism $\Phi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ given by $\Phi(f)=f|_{M_{\mathrm{prin}}}$. Note that $\Phi$ is injective. Assume that $\Phi(f)=\mathrm{Id}_{M_{\mathrm{prin}}}$. Since $M_{\mathrm{prin}}$ is dense and $M$ is Hausdorff, $f=\mathrm{Id}_{M}$ by Lemma 3.5. Furthermore, the map $\Phi$ is continuous and, therefore, smooth by [26, Corollary 3.50]. Hence, $\Phi$ is a Lie group homomorphism.
Fix $v\in T_{\mathrm{Id}_{M}}\operatorname{Isom}(M,\mathcal{F})$ and assume that $\Phi_{\ast_{\mathrm{Id}_{M}}}(v)=0$. We will show that $v=0$, which implies that $\Phi_{\ast_{\mathrm{Id}_{M}}}$ is injective. Since $\Phi$ is a Lie group homomorphism and $\Phi(\mathrm{Id}_{M})=\mathrm{Id}_{M_{\mathrm{prin}}}$, we have
(3.3) $\Phi(\exp_{\mathrm{Id}_{M}}(v))=\exp_{\mathrm{Id}_{M_{\mathrm{prin}}}}(\Phi_{\ast_{\mathrm{Id}_{M}}}(v))=\mathrm{Id}_{M_{\mathrm{prin}}}.$
By definition, $\Phi(\exp_{\mathrm{Id}_{M}}(v))=\exp_{\mathrm{Id}_{M}}(v)|_{M_{\mathrm{prin}}}$. Hence, equation (3.3) implies that the isometry $\exp_{\mathrm{Id}_{M}}(v)\colon M\to M$ restricted to the open and dense subset $M_{\mathrm{prin}}\subset M$ is the identity map, i.e., $\exp_{\mathrm{Id}_{M}}(v)|_{M_{\mathrm{prin}}}=\mathrm{Id}_{M_{\mathrm{prin}}}=\mathrm{Id}_{M}|_{M_{\mathrm{prin}}}$. Then, by Lemma 3.5, $\exp_{\mathrm{Id}_{M}}(v)=\mathrm{Id}_{M}$. This implies that $v=0$. Therefore, $\Phi_{\ast_{\mathrm{Id}_{M}}}$ is injective and it follows that
$\dim(\operatorname{Isom}(M,\mathcal{F}))\leq\dim(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})).$
Thus, to prove inequality (a), it suffices to show that
(3.4) $\dim(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
The projection $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is a fiber bundle with $(n-k)$-dimensional fibers (see [34, Theorem A] and [10, p. 41]). Hence, $M_{\mathrm{prin}}/\mathcal{F}\subset M/\mathcal{F}$ is a $k$-dimensional smooth manifold and inherits a Riemannian metric from $M_{\mathrm{prin}}$, making the map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ a Riemannian submersion. By the classical Myers–Steenrod theorem [37], both $\operatorname{Isom}(M_{\mathrm{prin}})$ and $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ are Lie groups. Moreover, by [28, II Theorem 3.1],
(3.5) $\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}))\leq\frac{k(k+1)}{2}.$
Consider now the group homomorphism $\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{\mathrm{Homeo}}(M_{\mathrm{prin}}/\mathcal{F})$ given by
$\Psi(f)(p^{\ast})=(\pi\circ f)(q),$
where $q\in\pi^{-1}(p^{\ast})$. Let $G=\ker(\Psi)$. Since the leaves of $\mathcal{F}$ are closed, by [10] and [34], for any principal leaf $L\in\mathcal{F}$ which has dimension $(n-k)$, a sufficiently small tubular neighborhood $\mathrm{Tub}(L)$ of $L$ is foliated diffeomorphic to
$\mathrm{Tub}(L)\cong\left(L\times D^{k},\{L\times\{x\}\mid x\in D^{k}\}\right),$
where $D^{k}$ denotes a $k$-dimensional disk. We point out that $\mathrm{Tub}(L)^{\ast}\cong D^{k}$. Thus, by definition, the group $G$ induces a trivial action on $D^{k}$, and $G$ acts effectively on the tubular neighborhood; otherwise, the group $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ would not act effectively. In this way, we get an effective action of $G$ on $L$. Moreover, this action is by isometries with respect to the induced metric of $M$ on $L$. Thus, by [28, II Theorem 3.1],
(3.6) $\dim(G)\leqslant\frac{(n-k)(n-k+1)}{2}.$
Since $\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is a continuous Lie group homomorphism, $\Psi$ is smooth. Moreover, $\Psi(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F}))$ is a Lie subgroup of $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$. Then, by the rank-nullity theorem and inequalities (3.5) and (3.6),
$\dim(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F}))$ $=\dim(G)+\dim(\Psi(\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})))$
$\leqslant\dim(G)+\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}))$
$\leqslant\frac{(n-k)(n-k+1)}{2}+\frac{k(k+1)}{2}.$
This last series of inequalities verifies inequality (3.4), which in turn implies inequality (a). ∎
Remark 3.6.
Observe that $f\in\ker(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast}))$ if and only if $f(L)\subset L$ for any leaf $L\in\mathcal{F}$. Now, for the Lie group homomorphism $\Phi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ in the proof of Theorem A (ii), if $f\in\ker\left(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast})\right)$, then $\Phi(f)(L)=L$ for any leaf $L\subset M_{\mathrm{prin}}$. This implies that $\Phi(f)\in\ker(\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}^{\ast}))$. Thus, we conclude that
$\Phi(\ker(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast}))\subset\ker(\Psi\colon\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}^{\ast})).$
This, together with the fact that $\Phi_{\ast_{\mathrm{Id}_{M}}}$ is injective, implies that
@@
$\dim(\ker(\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M^{\ast})))\leq\frac{k(k+1)}{2}.$
With this observation in hand, we may now finish the proof of Theorem A.
3.3. Proof of Theorem A (iii)
Consider the Lie group homomorphism $\Psi\colon\operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M/\mathcal{F})$. Since $\operatorname{Isom}(M,\mathcal{F})$ has maximal dimension, Remark 3.6 implies that the kernel $G=\ker\Psi$ has dimension equal to $k(k+1)/2$. Moreover, since $G$ acts effectively by isometries on any leaf, we conclude, by [28, II Theorem 3.1], that all the leaves are isometric to a round sphere $S^{k}$ or a round real projective space $\mathbb{R}P^{k}$. This implies that the foliation $\mathcal{F}$ is regular.
Equality in the dimension of $\operatorname{Isom}(M,\mathcal{F})$ implies that the dimension of $\operatorname{Isom}(M/\mathcal{F})$ is equal to $(n-k)(n-k+1)/2$. The proof of Theorem A (ii) in turn implies that $\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})=(n-k)(n-k+1)/2$. Since $M_{\mathrm{prin}}/\mathcal{F}$ is a Riemannian manifold, we conclude, by [28, II Theorem 3.1], that $M_{\mathrm{prin}}/\mathcal{F}$ is isometric to one of the following: an $(n-k)$-dimensional Euclidean space, an $(n-k)$-dimensional hyperbolic space with constant negative curvature, an $(n-k)$-dimensional round real projective space, or an $(n-k)$-dimensional round sphere. Therefore, $M_{\mathrm{prin}}/\mathcal{F}$ is complete. Since $M^{\ast}$ is the metric completion of $M_{\mathrm{prin}}/\mathcal{F}$, it follows that $M^{\ast}=M_{\mathrm{prin}}/\mathcal{F}$. Recall that over the principal leaves of the singular Riemannian foliation $\mathcal{F}$, the map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is a smooth Riemannian submersion. This implies that $\pi\colon M\to M/\mathcal{F}$ is a fiber bundle, and thus the desired conclusion follows. ∎
4. Proof of Theorem B
We will prove each item in Theorem B separately. Throughout this section, we let $\pi\colon X\to Y$ be a proper submetry between Alexandrov spaces and set =$\mathcal{F}=\{\pi^{-1}(y)\mid y\in Y\}$.
4.1. Proof of Theorem B (i)
Since $Y$ is an Alexandrov space, Theorem 2.18 (1) implies that $\operatorname{Isom}(Y)$ is a closed Lie group. Then, by Corollary 3.2, $\operatorname{Isom}(X,\mathcal{F})$ is a Lie group. Moreover, when $X$ is compact, $Y$ is also compact and, hence, $\operatorname{Isom}(X)$ and $\operatorname{Isom}(Y)$ are compact, and thus closed. Consider the map $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)(y)=(\pi\circ f)(x)$ for any $x\in\pi^{-1}(y)$ (cf. Section 2.7). By Proposition 3.1, $\operatorname{Isom}(X,\mathcal{F})$ is closed in $\operatorname{Isom}(X)$, since $\operatorname{Isom}(Y)$ is closed. Thus, we conclude that $\operatorname{Isom}(X,\mathcal{F})$ is compact. ∎
4.2. Proof of Theorem B (ii)
We now compute the upper bound on the dimension of the Lie group $\operatorname{Isom}(X,\mathcal{F})$. Suppose that $X$ has dimension $n\geq 1$ and $Y$ has dimension $0\leqslant k\leqslant n$, so that the regular leaves of the foliation $\mathcal{F}$ have dimension $n-k$. We will show that
(4.1) $\dim(\operatorname{Isom}(X,\mathcal{F}))\leqslant\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
Recall that a point $y\in Y$ is regular with respect to $\pi$ if $\pi$ is regular at each $x\in\pi^{-1}(y)$ (see [29, Definition 6.5]). The set of regular points $Y_{\mathrm{reg}}=\{y\in Y\mid y\mbox{ is regular with respec to }\pi\}$ contains the open and dense subset of thick points (see [29, End of Section 6]). Consequently, the set $X_{\mathrm{reg}}=\{x\in X\mid x\mbox{ is regular}\}$ contains an open and dense subset. Let $R_{X}=\{x\in X\mid\Sigma_{x}X\mbox{ is isometric to the unit round }S^{n-1}\}$. By Theorem 2.12, $R_{X}$ is dense. Hence, $R_{X}\cap X_{\mathrm{reg}}\neq\varnothing$.
Fix $\bar{x}\in R_{X}\cap X_{\mathrm{reg}}$. We may choose $\bar{x}$ so that $\Sigma_{\pi(\bar{x})}Y$ is isometric to the unit round $S^{k}$, since the set of points in $Y$ whose space of directions is isometric to the unit round $S^{k}$ is dense in $Y$ and $Y_{\mathrm{reg}}$ contains an open subset. Since $\bar{x}$ is regular, $d_{\bar{x}}\pi|_{H_{\bar{x}}}\colon H_{\bar{x}}\to\Sigma_{\pi(\bar{x})}Y$ is an isometry (see Section 2.4). Moreover, since $\Sigma_{\pi(\bar{x})}Y$ is round, there exists an isometry between $\Sigma_{\bar{x}}X$ and the spherical join $H_{\bar{x}}\star V_{\bar{x}}$. Since $\Sigma_{\bar{x}}X$ is isometric to the unit round $S^{n-1}$ and $H_{\bar{x}}$ is isometric to the unit round $S^{k-1}$, we conclude that $V_{\bar{x}}$ is isometric to the unit round $S^{n-k-1}$.
Consider now the Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)=\pi\circ f$ (cf. Section 2.7 and Section 3) and set $G=\ker(\Psi)$.
Given $f\in G$, we have $f(\bar{x})\in\pi^{-1}(\pi(\bar{x}))$, allowing us to identify the differential $d_{\bar{x}}f\colon\Sigma_{\bar{x}}X\to\Sigma_{f(\bar{x})}X$ (wich is an isometry) with an isometry
$(d_{\bar{x}}f)_{1}\star(d_{\bar{x}}f)_{2}\colon H_{\bar{x}}\star V_{\bar{x}}\to H_{f(\bar{x})}\star V_{f(\bar{x})}.$
Since $G$ acts on $\pi^{-1}(\pi(\bar{x}))$, we may consider the isotropy group $G_{\bar{x}}$. For $f\in G_{\bar{x}}$, we have then that $(d_{\bar{x}}f)_{1}=\mathrm{Id}_{H_{\bar{x}}}$ and $(d_{\bar{x}}f)_{2}\colon V_{\bar{x}}\to V_{\bar{x}}$ is an isometry. Thus, $G_{\bar{x}}$ acts by isometries on the unit round $S^{n-k-1}$. Moreover, this action is effective. Indeed, if $(d_{\bar{x}}f)_{2}=\mathrm{Id}_{V_{\bar{x}}}$, then $d_{\bar{x}}f=\mathrm{Id}_{\Sigma_{\bar{x}}X}$. Since $f(\bar{x})=\bar{x}$ and $f$ is an isometry of the Alexandrov space $X$, then [18, Lemma 3.2] implies that $f=\mathrm{Id}_{X}$. Hence, by the Myers–Steenrod theorem,
$\dim(G_{\bar{x}})\leq\frac{(n-k-1)(n-k)}{2}.$
Now we consider the orbit $G(\bar{x})\subset\pi^{-1}(\pi(\bar{x}))$. By [29, Korollar 7.5], the Hausdorff dimension of $\pi^{-1}(\pi(\bar{x}))$ is $n-k$. Thus, we conclude that the dimension of $G(\bar{x})=G/G_{\bar{x}}$ is bounded above by $n-k$. It follows then that
(4.2) $\begin{split}\dim(G)=\dim(G(\bar{x}))+\dim(G_{\bar{x}})&\leq n-k+\frac{(n-k-1)(n-k)}{2}\\[5.69046pt] &=\frac{(n-k+1)(n-k)}{2}.\end{split}$
(4.3) $\dim(\Psi(\operatorname{Isom}(X,\mathcal{F})))\leqslant k(k+1)/2.$
Then, by the first group isomorphism theorem, inequalities (4.2) and (4.3), and recalling that $G=\ker\Psi$,
(4.4) $\begin{split}\dim\left(\operatorname{Isom}(X,\mathcal{F})\right)&=\dim\left(G\right)+\dim\left(\Psi(\operatorname{Isom}(X,\mathcal{F})\right))\\[5.69046pt] &\leq\frac{(n-k)(n-k+1)}{2}+\frac{k(k+1)}{2}.\end{split}$
4.3. Proof of Theorem B (iii)
We now examine the case where equality in inequality (4.4) holds. We first determine the topology of the base space $Y$ of the submetry $\pi\colon X\to Y$.
Proposition 4.1.
Let $\pi\colon X\to Y$ be a proper submetry from an Alexandrov space $X$ of dimension $n$ onto an Alexandrov space $Y$ of dimension $0\leqslant k\leqslant n$. If
then $Y$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
@@ 4.3. Proof of Theorem B (iii)
Since $Y$ is $k$-dimensional, Theorem 2.18 implies that $\dim(\operatorname{Isom}(Y))\leq k(k+1)/2$. Since $\dim(\operatorname{Isom}(X,\mathcal{F}))$ is maximal, it follows from the proof of Theorem B (ii) that the subgroup $\Psi(\operatorname{Isom}(X,\mathcal{F}))\subset\operatorname{Isom}(Y)$ has dimension $k(k+1)/2$. Thus, $\operatorname{Isom}(Y)$ has dimension $k(k+1)/2$ and the conclusion follows from Theorem 2.18. ∎
We now determine the topology of the leaves and show that the submetry is a topological fibration, finishing the proof of item (iii) in Theorem  B.
@@
Let $\pi\colon X\to Y$ be a proper submetry from an Alexandrov space $X$ of dimension $n$ onto an Alexandrov space $Y$ of dimension $0\leqslant k\leqslant n$. If
then there exists a Lie group $G$ acting by isometries on $X$ such that the submetry $\pi$ is the composition of the submetry $\pi_{G}\colon X\to X/G$ and a submetry $\pi_{D}\colon X/G\to Y$ with discrete fibers. Moreover, all $G$-orbits are homeomorphic, and each leaf of $\pi$ is homeomorphic to a finite union of orbits of the $G$-action.
@@
Consider the group Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ given by $\Psi(f)=\pi\circ f$ and set $G=\operatorname{Ker}(\Psi)$. As stated in the proof of Theorem B (ii), we may consider $\bar{x}\in X$ such that $\Sigma_{\bar{x}}(X)$ is isometric to the unit round $S^{n-1}$ and $\Sigma_{\pi(\bar{x})}(Y)$ is isometric to a unit round $S^{k-1}$.
From our hypotheses and the proof of Theorem B (ii), we observe that the orbit $G(\bar{x})$ has Hausdorff dimenstion equal to $n-k$, which is the Hausdorff dimension of the leaf $\pi^{-1}(\pi(\bar{x}))$.
We consider the space $X/G$ endowed with the orbit projection distance $d^{\ast}$. Note that we have a submetry $\pi_{G}\colon(X,d_{X})\to(X/G,d^{\ast})$. Furthermore, since the orbits of $G$ are contained in the leaves of $\pi$ (i.e., $G(x)\subset\pi^{-1}(\pi(x))$ for any $x\in X$), we obtain a well defined submetry $\pi_{D}\colon(X/G,d^{\ast})\to(Y,d_{Y})$ such that $\pi=\pi_{D}\circ\pi_{G}$. Note also that the Hausdorff dimension of the orbits $G(x)$ is bounded by the Hausdorff dimension of the leaves, which is bounded by $n-k$. Moreover, the leaf $L_{\bar{x}}=\pi^{-1}(\pi(\bar{x}))$ has Hausdorff dimension equal to $n-k$. From our hypotheses and the proof of Theorem B (ii), the dimension of the orbit $G(\bar{x})$ must also be $n-k$. Hence, the orbit space $X/G$ has Hausdorff dimension $n-(n-k)=k$. This implies that the submetry $\pi_{D}$ has discrete fibers, and thus $\pi_{D}^{-1}(\pi(x))=\sqcup_{i\in\mathbb{N}}\{x^{\ast}_{i}\}$. Since $\pi=\pi_{D}\circ\pi_{G}$, it follows that
$L_{\bar{x}}=\bigsqcup_{i\in\mathbb{N}}G(x_{i}).$
By [29, Theorem 7.2], the connected components of each fiber of $\pi$ have positive distance to one another.
Note now that, by Proposition 4.1, the space $Y$ is a Riemannian manifold. Thus, every point $y\in Y$ is a regular point of both $\pi$ and $\pi_{G}$. From the definition of regular point we conclude that given $x^{\ast}\in X/G$, the set of horizontal directions $H_{x^{\ast}}(\pi_{D})$ of $\pi_{D}$ is isometric to $\Sigma_{\pi_{G}(x^{\ast})}Y=S^{k-1}$. Moreover, since the fibers of $\pi_{D}$ is discrete, $H_{x^{\ast}}(\pi_{D})=\Sigma_{x^{\ast}}(X/G)$. This implies that all leaves of $\pi$ and $\pi_{G}$ are regular leaves. By Corollary 2.17, each leaf of $\pi$ is homeomorphic to $L_{\bar{x}}$, and each $G$-orbit is homeomomrphic to $G(\bar{x})$.
Let $U\subset G$ be an open neighborhood of the identity element $e\in G$ and set $U_{i}=\{gx_{i}\mid g\in U\}$, $i\in\mathbb{N}$. Then we have that $\{gU_{i}\mid g\in G,\>i\in\mathbb{N}\}$ is an open cover of $L_{\bar{x}}$. Moreover, by construction, $gU_{i}\cap hU_{j}=\varnothing$ for all $g,h\in G$ and $i\neq j$. Since $L_{\bar{x}}$ is compact, there exists a finite subcover of $\{gU_{i}\mid g\in G,\>i\in\mathbb{N}\}$. It follows then that $L_{\bar{x}}$ must be the finite union of disjoint orbits $G(x_{1}),\cdots,G(x_{N})$. ∎
Proposition 4.3.
Let $\pi\colon X\to Y$ be a proper submetry from an Alexandrov space $X$ of dimension $n$ onto an Alexandrov space $Y$ of dimension $0\leqslant k\leqslant n$. If
then $(X,\mathcal{F})$ is foliated-homeomorphic to a fiber bundle over $Y$ whose fibers are isometric to a finite disjoint union of round spheres or a finite disjoint union of round projective spaces.

### v3

@@ 4.3 Proof of Theorem B(iii)(iii)
As stated in the proof of Theorem B(ii)(ii), we may fix $\bar{x}\in X$ such that $\Sigma_{\bar{x}}(X)$ is isometric to a unit round $S^{n-1}$ and $\Sigma_{\pi(\bar{x})}(Y)$ is isometric to a unit round $S^{k-1}$. From our hypotheses and the proof of Theorem B(ii)(ii), we observe that $\dim_{H}(G(\bar{x}))=n-k=\dim_{H}(L_{\bar{x}})$, where $L_{\bar{x}}=\pi^{-1}(\pi(\bar{x}))$ is the fiber containing $\bar{x}$ and $\dim_{H}(\cdot)$ denotes Hausdorff dimension.
Endow the orbit space $X/G$ with the orbital distance $d^{\ast}$ and consider the submetry $\pi_{G}\colon(X,\allowbreak d_{X})\to(X/G,d^{\ast})$. Since the orbits of $G$ are contained in the fibers of $\pi$ (i.e., $G(x)\subset\pi^{-1}(\pi(x))$ for any $x\in X$), we obtain a well-defined map $\pi_{D}\colon(X/G,d^{\ast})\to(Y,d_{Y})$ with $\pi=\pi_{D}\circ\pi_{G}$. Since $\pi$ and $\pi_{G}$ are submetries, $\pi_{D}$ is also a submetry. Hence, $\dim_{H}(X/G)=n-(n-k)=k$. This implies that the submetry $\pi_{D}$ has discrete fibers, and thus $\pi_{D}^{-1}(\pi(x))=\bigsqcup_{i\in\mathbb{N}}\{x^{\ast}_{i}\}$. Since $\pi=\pi_{D}\circ\pi_{G}$}, it follows that $L_{\bar{x}}=\bigsqcup_{i\in\mathbb{N}}G(x_{i})$. By [36, Theorem 7.2], the connected components of each fiber of $\pi$ have positive distance to one another. By hypothesis, $L_{\bar{x}}$ is connected. Hence, for $G^{0}\subset G$, the connected component of the identity, we have that $G(\bar{x})=G^{0}(\bar{x})$, $\pi=\pi_{G}$, and $L_{\bar{x}}=G^{0}(\bar{x})$. Thus, $L_{\bar{x}}$ with the intrinsic metric is a compact homogeneous inner metric space.
Since $\operatorname{Isom}(X)$ is a Lie group, $G^{0}$ is a Lie subgroup. Hence, $L_{\bar{x}}=G^{0}(\bar{x})=G^{0}/\bigl(G^{0}\bigr)_{\bar{x}}$ is a manifold and, in particular, locally compact and locally contractible. We also have $\dim_{\mathrm{top}}L_{\bar{x}}=\dim_{H}L_{\bar{x}}=n-k$. By [9] (cf. [7, Theorem 13]), $(L_{\bar{x}},d_{I})$ is a homogeneous Finsler manifold. Moreover, at $\bar{x}$, the space of directions $\Sigma_{\bar{x}}L_{\bar{x}}$ is isometric to the unit round $S^{n-k-1}$ and the metric tangent cone is isometric to the (nk)(n-k)-dimensional Euclidean space. Since $G^{0}$ acts transitively and by isometries on $L_{\bar{x}}$, this holds at every point in $L_{\bar{x}}$. Therefore, the Finsler norm is Euclidean, and it follows that $(L_{\bar{x}},d_{I})$ is a homogeneous Riemannian manifold (see also [6, Theorem 7]).
Therefore, $L_{\bar{x}}$ is a union of Riemannian manifolds and, hence, a compact Riemannian manifold with finitely many connected components. By hypothesis, the isometry group of $L_{\bar{x}}$ has maximal dimension (nk)(nk+1)/2(n-k)(n-k+1)/2. Hence, by [33, Chapter II, Theorem 3.1], $L_{\bar{x}}$ is isometric to a round $S^{n-k}$ or a round $\mathbb{R}P^{n-k}$.
By Lemma 4.3, the space $Y$ is a Riemannian manifold. Thus, every point $y\in Y$ is a regular point of $\pi=\pi_{G}$. From the definition of regular point, we conclude that given $x^{\ast}\in X/G$, the set of horizontal directions $H_{x^{\ast}}(\pi_{D})$ of $\pi_{D}$ is isometric to $\Sigma_{\pi(x^{\ast})}Y=S^{k-1}$. Moreover, $H_{x^{\ast}}(\pi_{D})=\Sigma_{x^{\ast}}(X/G)$. This implies that all fibers of $\pi=\pi_{G}$ are regular. By Corollary 2.17, each fiber of $\pi$ is homeomorphic to $L_{\bar{x}}=G(\bar{x})$. That is, all the fibers are homogeneous spaces isometric to round spheres or round real projective spaces. ∎
Lemma 4.6.
@@
If the fibers of $\pi\colon X\to Y$ are connected, then $\operatorname{Isom}(X,\mathcal{F})$ acts transitively on $X$.
@@
By Theorem B(ii)(ii) and equality in (4.3), the image $\Psi(\operatorname{Isom}(X,\mathcal{F}))\subset\operatorname{Isom}(Y)$ has the same dimension as $\operatorname{Isom}(Y)$. Hence, $\Psi(\operatorname{Isom}(X,\mathcal{F}))$ contains $\operatorname{Isom}(Y)^{0}$, the identity component of $\operatorname{Isom}(Y)$. By Lemma 4.3, $Y$ is isometric to $\mathbb{R}^{k}$, $\mathbb{H}^{k}$, a round $S^{k}$ or a round $\mathbb{R}P^{k}$. Thus, $\operatorname{Isom}(Y)^{0}$ is isomorphic to $\mathbb{R}^{k}\rtimes\mathrm{SO}(k)$ (Euclidean space), $\mathrm{SO}(k,1)$ (hyperbolic space), or $\mathrm{SO}(k+1)$} (sphere and real projective space). In each case, $\operatorname{Isom}(Y)^{0}$ acts transitively on $Y$. Therefore, $\Psi(\operatorname{Isom}(X,\mathcal{F}))$ acts transitively on $Y$.
Let $x,\bar{x}\in X$ and choose $h\in\operatorname{Isom}(X,\mathcal{F})$ with $\Psi(h)(\pi(x))=\pi(\bar{x})$. Then, by the definition of $\Psi$, we have $h(x)\in\pi^{-1}(\pi(\bar{x}))$. Since $G\subset\operatorname{Isom}(X,\mathcal{F})$ acts transitively on $\pi^{-1}(\pi(\bar{x}))$ by the proof of Lemma 4.5, there exists $g\in G=\ker(\Psi)$ with $g(h(x))=\bar{x}$. Thus, we conclude that $\operatorname{Isom}(X,\mathcal{F})$ acts transitively on $X$. ∎
@@
If all fibers of $\pi\colon X\to Y$ are connected, then $X$ and $Y$ are Riemannian manifolds, and $\pi\colon X\to Y$ is a smooth Riemannian submersion.
@@
By Lemma 4.6, $X$ is a homogeneous Alexandrov space. Hence, by [6, Theorem 7], $X$ is isometric to a homogeneous Riemannian manifold $(M,g)$. By Lemma 4.5, each fiber is an orbit $G^{0}(x)$ of $G=\ker(\Psi)\subset\operatorname{Isom}(X,\mathcal{F})$. Since $X$ is now Riemannian and $G\subset\operatorname{Isom}(X)$ acts by isometries, each fiber is a smooth embedded submanifold of $X$. Finally, by [39, Lemma 13.1] implies that $\pi\colon X\to Y$ is a smooth map. ∎
4.4 Disconnected fiber case
Suppose now that the fibers of $\pi\colon X\to Y$ have one or more connected components. Recall our standing assumption that the submetry $\pi\colon X\to Y$ is proper and equality in (4.3) holds. Recall that by Lemma 4.3, $Y$ is a $k$-dimensional Riemannian manifold isometric to Euclidean space, hyperbolic space, a round sphere, or a round real projective space. We distinguish two cases, depending on whether $Y$ is compact.
Lemma 4.8.
If the fibers of $\pi\colon X\to Y$ are possibly disconnected and $Y$ is isometric to Euclidean or hyperbolic space, then $X$ is homeomorphic to $Y\times F$ with $F$ either a round sphere or a round real projective space.